OpenStudy (anonymous):

Lastly, make "t" the subject of S = (1/2)(a)(t^2). (Consider only t > 0)

jhonyy9 (jhonyy9):

S=(1/2)(a)(t^2) --- S=(at^2)/2 /*2 2S=at^2 --- t^2 =(2S)/a --- t_1,_2=+/-sqrt(2S/a)

OpenStudy (anonymous):

Hey thanx !

jhonyy9 (jhonyy9):

was my pleasure